OpenStudy (anonymous):

1 3/5 divided by 5/8 = ?

OpenStudy (anonymous):

1 1 1/5 2 14/25 2 2/5

OpenStudy (alexandervonhumboldt2):

convert into improper fraction: \[a \frac{ b }{ c }=\frac{ a*c+b }{ c }\] \[1\frac{ 3 }{ 5 }=\frac{ 1*5+3 }{ 5 }=\frac{ 8 }{ 5 }\] now divide: \[\frac{ 8 }{ 5 }/\frac{ 5 }{ 8 }=\frac{ 8 }{ 5 }*\frac{ 8 }{ 5}=\frac{ 64 }{ 25 }=2\frac{ 14 }{ 25 }\]